# SimMerge: Learning to Select Merge Operators from Similarity Signals

### v1

@@ 2.3 Bandit View and Evaluation Protocol
The offline selector $f_{\mathrm{op}}$ works well under the distribution of model pairs and tasks seen during training, but it requires to be retrained when new models or tasks are introduced. To support such scenarios without retraining the similarity encoder, we also view operator selection as a contextual bandit problem and add a linear bandit layer on top of the learned features. We then evaluate both the offline and bandit variants in pairwise and multiway regimes.
Contextual bandit formulation. Each merge step, either a standalone 2-way merge or a step inside a multiway plan, defines a decision round. We observe a context vector $s\in\mathbb{R}^{m}$ derived from pre-merge similarity features. For pairwise merges, we take $s=\tilde{x}(m_{a},m_{b},t)$. For multiway merges, $s$ is constructed analogously but may include propagated similarity features for intermediate merges. We then choose an action $a$ from the finite set of operators $\mathcal{A}=\{\textsc{Linear}{},\textsc{Slerp}{},\textsc{Ties}{}\}$, apply the corresponding merge operator, and observe a scalar reward $r(a)$ given by the downstream evaluation score. Rewards for unchosen operators are not observed. The objective is to learn a policy $\pi(s)$ mapping contexts to operators that maximizes cumulative reward, or equivalently minimizes regret with respect to an oracle that always picks the best operator for each context.
We adopt a neural-linear design. An MLP feature map $g_{\phi}$ transforms the context $s$ into a representation $z(s)=g_{\phi}(s)$. We initialize $g_{\phi}$ and warm-start it using the logged pairwise data described below, then keep it fixed during online adaptation. On top of $z(s)$ we fit a linear contextual bandit: for each operator $a\in\mathcal{A}$ we assume a linear reward model
@@
with an unknown parameter vector $w_{a}$ and a Gaussian posterior $\mathcal{N}(\hat{w}_{a},\Sigma_{a})$ maintained via Bayesian linear regression. We consider both LinUCB and linear Thompson sampling (LinTS) [Abbasi-Yadkori et al., 2011; Agrawal & Goyal, 2013]. Empirically LinTS produces lower regret and higher downstream performance in our setting, and we therefore use it as the main bandit variant. Since we have a small number of operators and a low-dimensional representation, posterior updates can be done with rank-one updates in $O(d^{2})$ time per round, where $d$ is the dimension of $z(s)$.
Warm-start and online adaptation. The bandit is initialized using the fully-observed pairwise merge dataset constructed in the offline setting. For each historical 2-way merge we know the context and the realized utilities $s_{o}$ for all operators $o\in\mathcal{A}$, which corresponds to full-information feedback. These logs are used once to fit a Bayesian linear model for each arm, yielding initial posteriors $(\hat{w}_{a},\Sigma_{a})$. This warm-start phase anchors the reward model before any online interaction and reduces the amount of exploration required when new tasks or models are introduced.
After warm-start, we introduce a distribution shift by adding a checkpoint trained on a different task that did not appear in the logged data. For each new merge involving this checkpoint, we compute the context $s$ from pre-merge similarity signals, query the bandit to select an operator $a$, execute $M_{a}$, and observe the resulting reward $r(a)$ from downstream evaluation. Only the posterior of arm $a$ is then updated. This is the partial-feedback regime where counterfactual rewards for unchosen operators are not observed.
@@ 3 Experimental Setup
We evaluate SimMerge on four domains: code generation, mathematical reasoning, multilingual understanding and RAG. Within each model scale, all checkpoints share the same pretrained base and architecture, so performance differences arise from task-specific fine-tuning and merging rather than capacity mismatches. At the 7B scale, we use Command-A 7B as the shared pretrained base for all domain-specialized checkpoints. At the larger scale, we use Command-A  [^1] B as the shared base for all  [^1] B experiments [Cohere et al., 2025].
Given a target task $t$ and a set of $k$ models to merge, we compare three fixed merge operators, Linear, SLERP, and TIES, against our learned selector SimMerge. The fixed baselines apply their respective operator at every merge step within a configuration. For $k>2$, they also use a fixed merge order. In contrast, SimMerge does not introduce a new merge rule. It selects among {Linear, SLERP, TIES} from pre-merge similarity signals (Section 2) and applies the selected operator once per merge step. For $k>2$, SimMerge additionally selects a favorable merge order using the same similarity signals. Any improvement of SimMerge over the fixed baselines therefore comes from predicting the operator choice and, in the multiway setting, the merge order, without running a merge-and-evaluate search.
@@ 3.1 Experts and Auxiliary Models
Tasks and experts. For each of the four domains: code, math, multilingual, RAG, we fine-tune a shared base model on task-specific data to obtain a task expert checkpoint $m^{\text{exp}}_{t}$. This expert provides a strong reference point for task $t$, since it is the best single model we obtain by standard fine-tuning on $t$.
Auxiliary models. We refer to any model that is not fine-tuned on $t$ as an auxiliary (off-domain) model for task $t$. Concretely, when evaluating task $t$, the experts of the other domains act as auxiliaries. In the simplest case, we merge the task expert $m^{\text{exp}}_{t}$ with a single auxiliary $m^{\text{aux}}_{t}$. For instance, when evaluating on the code task, the code expert is $m^{\text{exp}}_{\text{code}}$ and the math expert is treated as an auxiliary $m^{\text{aux}}_{\text{code}}$. In larger-$k$ merges, we combine one expert with multiple auxiliaries $\{m^{\text{aux}}_{t,j}\}_{j}$. We always report the standalone performance of $m^{\text{exp}}_{t}$ and of the corresponding auxiliary model(s) for $t$ as natural upper- and lower-bound baselines.
Across all offline 7B experiments we use $85$ unique domain-specialized checkpoints: $23$ code, $24$ math, $24$ multilingual, and $15$ RAG models. For the online bandit experiments we introduce an additional $15$ instruct checkpoints at 7B, for a total of $100$ distinct 7B models. At the larger  [^1] B scale, we use additional $18$ task-specific checkpoints (5 code, 5 math, 4 multilingual, and 4 RAG), used only for evaluation.
@@ 3.3 Evaluations
Per-task evaluations. For each task $t$, we evaluate on a small collection of standard benchmarks (details in Appendix E). Math is evaluated on MATH [Hendrycks et al., 2021] and GSM8K [Zeng et al., 2023] using exact match. Multilingual performance is measured on an internal multilingual QA suite together with MGSM [Shi et al., 2022], using win-rates and accuracy. Code generation is evaluated with pass@1 on HumanEval_Python [Chen, 2021] and MBPP+ benchmark [Liu et al., 2023]. For RAG, we use TauBench [Yao et al., 2025] and BFCL [Patil et al., 2025] with accuracy and F1. In the instruction-following setting used for the bandit experiments, we report scores on IFEval [Zhou et al., 2023]. We run each evaluation three times with different seeds and report the mean score.
Task-level aggregation. To obtain a single score per task, we average the metric over all evaluations associated with that task. All aggregation is done separately for each merging method, so that comparisons between experts, auxiliaries, and merges are on the same task-specific scale.
@@ 3.4 Metrics
We evaluate merge quality using both absolute task performance and normalized improvements over two natural baselines. For a model $m$ on task $t$, define the expert baseline score
@@
and the auxiliary baseline score $s_{\text{aux}}(t)$ as follows: for pairwise merges, $s_{\text{aux}}(t):=\mathrm{score}(\text{aux}_{t},t)$ where $\text{aux}_{t}$ is the single auxiliary model paired with $\text{expert}_{t}$ and for multiway merges,
@@
where $\mathcal{A}_{t}$ is the set of auxiliary models included in that merge configuration. We then report normalized improvements:
@@
We aggregate over tasks by macro-averaging $\Delta_{\text{expert}}(m,t)$ and $\Delta_{\text{aux}}(m,t)$ across $t\in\mathcal{T}$. These normalized metrics account for scale differences across tasks and benchmarks and enable fair cross-domain comparisons. In addition, we report absolute performance (macro-averaged $\mathrm{score}(m,t)$) to anchor relative improvements on the original task scales.
In addition to $\Delta_{\text{expert}}$ and $\Delta_{\text{aux}}$, we also report results using the gap closed metric, which summarizes how much of the performance difference between the auxiliary baseline and the task expert is recovered by a merged model. For a model $m$ and task $t$, we define
@@
Under this normalization, $0\%$ corresponds to auxiliary performance and $100\%$ corresponds to expert performance. Values above $100\%$ indicate exceeding the expert, while negative values indicate performance below the auxiliary baseline.
@@ 3.5 Selector Architecture and Training Details
Offline classifier. Our offline selector is a lightweight MLP that takes the similarity features $\tilde{x}(m_{a},m_{b},t)$ as input and outputs a categorical distribution over $\{\text{Linear},\text{SLERP},\text{TIES}\}$. We use a two-layer network with ReLU activations and Adam optimization. We tune basic hyperparameters (hidden width, learning rate, batch size, dropout) on a held-out validation subset of the pairwise merge dataset and use early stopping on validation accuracy.
Bandit-based selector. For the online experiments we adopt the neural-linear contextual bandit described in Section 2.3. We instantiate the bandit feature map as a separate MLP encoder $g_{\phi}$, and take z(x~)z($\tilde${x}) to be its final hidden-layer representation. We warm-start this encoder and the per-operator Bayesian linear models using the training pairwise merge data described in Section 2.3, and then keep the encoder fixed during online adaptation. On top of z(x~)z($\tilde${x}), we maintain a Bayesian linear model for each operator and use linear Thompson sampling [Abbasi-Yadkori et al., 2011; Agrawal & Goyal, 2013] as our primary bandit policy, tuning its exploration and regularization hyperparameters on a validation split of the logged data. This policy is then used to select operators when new tasks or models, for instance, a new set of instruct checkpoints are introduced, without retraining the encoder.
@@ 4 Results
Across code, math, multilingual, and RAG, predictive merge selection consistently improves the trade-off between retaining expert performance and recovering off-domain capability. Section 4.1 shows that in pairwise expert-auxiliary merges, SimMerge achieves the strongest overall balance, reducing degradation relative to the expert while improving over auxiliary baselines. Section 4.2 shows that these gains extend to 3-way and 4-way merges, where SimMerge continues to outperform fixed operators on both expert-retention and auxiliary-recovery metrics. Because merge operators are not associative, merge order is an additional degree of freedom. Section 4.3 shows that the ordering selected by SimMerge yields consistent improvements over random orders. Section 4.4 shows that the same selector trained at 7B transfers to a  [^1] B model without retraining and again achieves the best expert/auxiliary trade-off. Finally, Section 4.5 considers an online setting with distribution shift and shows that a bandit variant learns to select operators efficiently and closely approaches oracle selection.
@@ 4.1 Learning to Choose the Right Merge Operator in the Pairwise Setting
We begin by examining the fundamental question underlying merge selection: given two fine-tuned models, can we predict which operator will produce the strongest composite model? Pairwise (2-way) merges offer the cleanest setting to answer this question and form the basis for training and validating SimMerge. For each task expert and auxiliary checkpoint, we compute pre-merge similarity features, allow SimMerge to choose among Linear, Slerp, and Ties, and compare the resulting merged model against the three fixed operators. This design isolates the contribution of operator selection, rather than changes in merging mechanics.
@@
Figure 2: Fraction of the expert–auxiliary performance gap closed by each merge method across Code, Math, Multilingual, and RAG tasks. SimMerge consistently recovers a larger fraction of expert performance than fixed merge operators across all domains.
@@
Figure 2 reports the fraction of the expert-auxiliary performance gap closed by each merge method across Code, Math, Multilingual, and RAG. Overall, the fixed operators exhibit substantial task dependence: an operator that performs well on one domain can be weak, or even harmful on another. For example, Linear is competitive on Code (69.0%) but slightly underperforms the auxiliary baseline on RAG (-0.4%), while TIES is strong on Math (83.3%) yet regresses below the auxiliary baseline on Multilingual (-10.2%). This variability highlights that there is no universally dominant fixed merge rule across domains.
In contrast, SimMerge consistently closes the largest fraction of the expert gap in all four domains. On Code, SimMerge closes 84.2% of the gap, outperforming the strongest fixed baseline (Linear at 69.0%) by +15.2 percentage points. On Math, SimMerge achieves 94.3% gap closed, exceeding all fixed operators (83.3% for TIES, 72.2% for SLERP, 61.5% for Linear), indicating that similarity-driven operator selection can recover near-expert performance in settings where multiple operators are viable but none is reliably optimal.
The Multilingual and RAG results further emphasize robustness. Multilingual is the most challenging domain for fixed operators: Linear closes 37.1% of the gap, SLERP closes 12.9%, and TIES drops below the auxiliary baseline (-10.2%). SimMerge reaches 46.9% gap closed, improving over the best fixed baseline (Linear) by +9.8 points and avoiding the severe regressions seen with TIES. On RAG, the gap closed values are lower overall, suggesting a larger remaining discrepancy between merged and expert performance. Still, SimMerge achieves the strongest result (34.8%), outperforming SLERP (23.6%) and TIES (6.4%), while Linear slightly underperforms the auxiliary (-0.4%). Macro-averaged across the four domains in Figure 2, SimMerge closes 65.0% of the expert-auxiliary gap, compared to 41.8% for the best single fixed operator (Linear); Appendix F.4 reports the calculation. Taken together, these patterns suggest that SimMerge not only improves peak performance where fixed operators already work well, but also provides a meaningful robustness advantage on domains where operator choice is harder.
Although SimMerge selects only among Linear, Slerp, and Ties, it can outperform all three fixed baselines because the fixed operators apply a single rule uniformly across merges, whereas SimMerge selects the operator per instance. Different expert-auxiliary pairs exhibit varying structural and functional relationships, including parameter alignment, scale differences, sign- and magnitude-based conflict patterns, functional similarity on probe inputs, and overall parameter geometry that strongly influence which operator succeeds. No single operator handles all such conditions well; however, these properties are reflected in the similarity features observed by SimMerge, enabling it to adapt its selection to each pair and approximate the per-instance optimal merge.
Overall, these results show that predictive operator selection based on similarity signals makes merging substantially more reliable than using a single fixed rule. By choosing the operator per expert-auxiliary pair, SimMerge avoids large regressions that occur when the wrong operator is applied and consistently recovers a larger fraction of expert performance across domains.
@@
Figure 3: Per-task percentage change in performance for each merge method. Blue markers show $\Delta_{\text{aux}}$ (change vs. auxiliary; higher is better) and red markers show $\Delta_{\text{expert}}$ (change vs. task expert; closer to 0 indicates less degradation).
Figure 3 makes this explicit by plotting the normalized percentage change relative to the auxiliary (blue) and the task expert (red). Across all tasks, SimMerge achieves the largest gains over the auxiliary model while incurring the smallest degradation relative to the expert. For example, it improves by +29.4+29.4% on Math and +33.8+33.8% on RAG while degrading by only 1.4-1.4% on Math and 2.4-2.4% on Code. In contrast, fixed operators show clear trade-offs: Linear can approach the expert on some tasks but fails to improve over the auxiliary on RAG, Slerp performs well on Math but weakly elsewhere and Ties even underperforms the auxiliary on Multilingual.
@@
These pairwise results show that similarity-driven operator selection reliably chooses a near-optimal merge rule for each expert-auxiliary pair, producing merged models that consistently improve over the auxiliary baseline while remaining closest to the task expert across all four domains. Appendix G analyzes how similarity signals correlate with outcomes and highlights the performance tails where fixed operators exhibit large regressions, helping explain why per-instance operator selection is critical for robust merging.

### v2

@@ 2.4 Bandit View and Evaluation Protocol
The offline selector performs well on the distribution of model pairs and tasks seen during training, but it must be retrained when new checkpoints, tasks, or operators are introduced. To support such scenarios without retraining the similarity feature computation, we cast operator selection as a contextual bandit and add a neural-linear bandit layer on top of the precomputed similarity features. As new model pairs and tasks appear, we compute their contexts from the same fixed feature pipeline and update the bandit online using the newly observed rewards.
Contextual bandit formulation. Each merge step, either a standalone two-way merge or a step within a multiway plan, defines a decision round. We observe a context vector $s\in\mathbb{R}^{m+d_{c}}$ derived from pre-merge similarity features. For pairwise merges, $s=\tilde{x}(m_{a},m_{b},t)$. For multiway merges, $s$ is constructed analogously and includes propagated similarity features for intermediate merges. We then choose an action $a$ from the operator set $\mathcal{O}$, apply the corresponding merge operator, and observe a scalar reward $r(a)$ given by downstream evaluation. Rewards for unchosen operators are not observed. The objective is to learn a policy $\pi(s)$ that maximizes cumulative reward, or equivalently minimizing regret relative to an oracle that always selects the best operator for each context.
We adopt a neural-linear design. An MLP feature map $g_{\phi}$ transforms the context $s$ into a representation $z(s)=g_{\phi}(s)$. We warm-start $g_{\phi}$ using the logged pairwise data described below, then keep it fixed during online adaptation. On top of $z(s)$ we fit a linear contextual bandit: for each operator $a\in\mathcal{O}$ we assume a linear reward model
@@
with an unknown parameter vector $w_{a}$ and a Gaussian posterior $\mathcal{N}(\hat{w}_{a},\Sigma_{a})$ maintained via Bayesian linear regression. We consider both LinUCB and linear Thompson sampling (LinTS) [Abbasi-Yadkori et al., 2011; Agrawal & Goyal, 2013]. Empirically, LinTS produces lower regret and higher downstream performance in our setting, and we therefore use it as our main bandit variant. Since we have a small number of operators and a low-dimensional representation, posterior updates can be done with rank-one updates in $O(d^{2})$ time per round, where $d$ is the dimension of $z(s)$.
Warm-start and online adaptation. The bandit is initialized using the fully-observed pairwise merge dataset from the offline setting. For each historical 2-way merge we observe the context $s$ and the utilities $\mathdollar U(\mathsf{a}(m_{a},m_{b}),t)$ for all operators $a\in\mathcal{O}$, which corresponds to full-information feedback. This warm-start phase anchors the reward model before any online interaction and reduces the amount of exploration required when new tasks or models are introduced.
After warm-start, we introduce a distribution shift by adding a checkpoint trained on a different task that did not appear in the logged data. For each new merge involving this checkpoint, we compute the context $s$ from pre-merge similarity signals, query the bandit to select an operator $\mathsf{a}$, execute the merge, and observe the resulting reward $r(a)$. Only the posterior corresponding to the chosen arm $a$ is then updated. This is the partial-feedback regime where counterfactual rewards for unchosen operators are not observed.
@@ 3 Experimental Setup
We evaluate SimMerge on four domains: code generation, mathematical reasoning, multilingual understanding and RAG. All experts share a common pretrained base (Command-A 7B or  [^1] B [Cohere et al., 2025]), ensuring differences arise from fine-tuning and merging rather than capacity.
For a target task $t$ and a set of $k$ models, we compare three fixed merge operators – Linear, Slerp, and TIES against our learned selector, SimMerge. Fixed operators apply the same rule at every merge step, using a predetermined order when $k>2$. In contrast, SimMerge chooses an operator (and, for $k>2$, a merge order) based on pre-merge similarity features, allowing performance gains without exhaustive merge-and-evaluate search.
@@ 3.1 Experts and Auxiliary Models
For each domain $t$, we fine-tune a shared base model on task-specific data and designate the best resulting checkpoint as the task expert $m^{\text{exp}}_{t}$. When evaluating task $t$, any model not fine-tuned on $t$ is treated as an auxiliary model; thus, experts from other domains act as auxiliaries for $t$. In all merge configurations, we combine one task expert with one or more auxiliary models. We report the standalone performance of task experts and auxiliary models as reference upper/lower bound baselines for task $t$, noting the best-performing expert or auxiliary may vary across metrics or configurations.
Across offline experiments at 7B, we use 85 domain-specialized checkpoints: 23 code, 24 math, 24 multilingual, and 15 RAG. For online bandit experiments, we add 15 instruction-tuned 7B checkpoints, yielding 100 distinct 7B models in total. At the  [^1] B, we evaluate on 18 additional task-specific checkpoints (5 code, 5 math, 4 multilingual, and 4 RAG).
@@ 3.3 Evaluations
For each task $t$, we evaluate on a small collection of standard benchmarks: . Math is evaluated on MATH [Hendrycks et al., 2021] and GSM8K [Zeng et al., 2023]; code generation on HumanEval [Chen, 2021] and MBPP+ [Liu et al., 2023]; multilingual understanding on MGSM [Shi et al., 2022] and an internal multilingual QA suite; retrieval-augmented generation (RAG) on TauBench [Yao et al., 2025] and BFCL [Patil et al., 2025]; and instruction-following on IFEval [Zhou et al., 2023] (details in Appendix E). Metrics follow standard task conventions; we repeat each evaluation three times with different seeds and report the mean. For each model and task $t$, we report a single task score given by the unweighted mean across $t$’s benchmarks.
@@ 3.4 Metrics
We evaluate merge quality using both absolute task performance and normalized percentage change relative to two natural baselines: (i) the task expert and (ii) the auxiliary model(s) involved in a merge. Reporting normalized change relative to these baselines enables fair comparison across tasks with different difficulty and score scales.
For a task $t$, we define the expert baseline score as
@@
where $\text{expert}_{t}$ is the model fine-tuned specifically for task $t$.
We define the auxiliary baseline score $s_{\text{aux}}(t)$ differently depending on the merge configuration. For pairwise merges,
$s_{\text{aux}}(t):=\mathrm{score}(\text{aux}_{t},t),$
where $\text{aux}_{t}$ is the single auxiliary model paired with the expert. For multiway merges, we define
@@
where $\mathcal{A}_{t}$ is the set of auxiliary models included in the merge configuration.
For a merged model $m$ evaluated on task $t$, we compute normalized percentage change relative to each baseline:
@@
These metrics quantify relative gains over the expert and auxiliary baselines, respectively, and account for differences in scale across tasks.
In addition, we report the gap closed metric, which measures how much of the performance gap between the auxiliary baseline and the task expert is recovered by the merged model:
@@
Under this normalization, $0\%$ corresponds to auxiliary performance and $100\%$ corresponds to expert performance. Values above $100\%$ indicate that the merged model exceeds the expert, while negative values indicate performance below the auxiliary baseline.
All normalized metrics are aggregated across tasks by macro-averaging over $t\in\mathcal{T}$. Absolute performance is also reported by macro-averaging the raw task scores to anchor normalized percentage change to the original evaluation scales.
@@ 3.5 Selector Architecture and Training Details
Offline classifier. We train a lightweight MLP that takes the similarity features $\tilde{x}(m_{a},m_{b},t)$ as input and outputs the predicted utility for the merge on task $t$. We use a two-layer network with ReLU activations and Adam optimization. We tune basic hyperparameters (hidden width, learning rate, batch size, dropout) on a held-out validation subset of the pairwise merge dataset and use early stopping on validation accuracy.
Bandit-based selector. We adopt the neural-linear contextual bandit described in Section 2.4. We instantiate the bandit feature map as a separate MLP encoder $g_{\phi}$, and take z(x~)z($\tilde${x}) to be its final hidden-layer representation. We warm-start this encoder and the per-operator Bayesian linear models using the training pairwise merge data described in Section 2.4, and then keep the encoder fixed during online adaptation. On top of z(x~)z($\tilde${x}), we maintain a Bayesian linear model for each operator and use linear Thompson sampling [Abbasi-Yadkori et al., 2011; Agrawal & Goyal, 2013] as our primary bandit policy, tuning its exploration and regularization hyperparameters on a validation split of the logged data. This policy is then used to select operators when new tasks or models, for instance, a new set of instruct checkpoints are introduced, without retraining the encoder.
@@ 4 Results
Across code, math, multilingual, and RAG, predictive merge selection with SimMerge consistently improves the performance of the final merged model. We show these gains for pairwise and multiway merges in Sections 4.1 and 4.2, respectively. In Section 4.3, we demonstrate the importance of merge-order prediction for SimMerge. Section 4.4 shows that a SimMerge selector trained on 7B pairwise merges transfers to  [^1] B 3-way merges without retraining. Finally, Section 4.5 considers an online setting with distribution shift and shows that a bandit variant learns to select operators efficiently, closely approaching oracle selection.
@@ 4.1 Learning to Choose the Right Merge Operator in the Pairwise Setting
We begin by examining the central question underlying merge selection: given two fine-tuned models, can we predict which operator will produce the best merged model? Pairwise (2-way) merges offer the cleanest setting to answer this question and form the basis for training and validating SimMerge.
For each expert-auxiliary checkpoint pair, we compute pre-merge similarity features and use SimMerge to select among Linear, Slerp, and TIES. We then compare the resulting merged model against each fixed operator. This design isolates the contribution of operator selection without changing other merging hyperparameters.
@@
Figure 2: Percentage of the expert–auxiliary performance gap closed by each merge method across Code, Math, Multilingual, and RAG tasks. SimMerge consistently recovers a larger fraction of expert performance than fixed merge operators across all domains.
@@
Figure 2 reports the fraction of the expert-auxiliary performance gap closed by each merge method across domains. Fixed operators exhibit substantial task dependence, with different operators excelling on different tasks and no single operator performing well across all domains. For example, Linear is competitive on Code (69.0%) but slightly underperforms the auxiliary baseline on RAG (-0.4%), while TIES is strong on Math (83.3%) yet regresses below the auxiliary baseline on Multilingual (-10.2%).
@@
Figure 3: Per-task normalized percentage change in performance for each merge method. Blue markers show $\Delta_{\text{aux}}$ and red markers show $\Delta_{\text{expert}}$ (closer to 0 indicates less degradation).
In contrast, SimMerge consistently closes the largest fraction of the expert gap in all four domains. It achieves 84.2% GapClosed on Code, 94.3% on Math, 46.9% on Multilingual, and 34.8% on RAG, outperforming the strongest fixed operator in each case. Macro-averaged across the four domains, SimMerge closes 65.0% of the expert-auxiliary gap, compared to 41.8% for the best single fixed operator (Linear).
The advantages of predictive operator selection is further illustrated in Figure 3, which shows the normalized percentage change in merged model performance relative to the auxiliary baseline (blue) and the task expert (red). Across all four domains, SimMerge achieves the largest gains over the auxiliary model while incurring the smallest degradation relative to the expert.
@@
These pairwise results indicate that similarity-driven operator consistently chooses an effective merge rule for each expert-auxiliary pair, producing merged models that improve over the auxiliary baseline while remaining closest to expert performance across all domains. Appendix G further analyzes how similarity signals correlate with merge outcomes and highlights the performance tails where fixed operators exhibit large regressions, underscoring why per-instance operator selection is critical for robust merging.